OpenStudy (anonymous):

I need help with this fast. Please help! Find the complementary event of the given event. a number less than 5 on a roll of a number cube with faces labeled 1– 6

OpenStudy (anonymous):

p=2/3

OpenStudy (anonymous):

A. rolling a 5 B. rolling a 5 or 6 C. rolling a 4 or 5 D. rolling a 4 Those are my choices

OpenStudy (epoweritheta):

b

OpenStudy (anonymous):

Thanks @epoweritheta

OpenStudy (anonymous):

Can you help me answer one more?

OpenStudy (epoweritheta):

ya

OpenStudy (anonymous):

Determine the sum of the given probabilities. Are the events complementary? P(red) = 2/5 ; P(blue) = 3/5

OpenStudy (epoweritheta):

give the full Q , otherwise we may not know that the events are disjoint

OpenStudy (anonymous):

that's the Question but the answers are, A. 1/2 the events are complimentary. B. 1/2 the events are not complimentary. C. 1 events are complimentary. D. 1 events are not complimentary.

OpenStudy (epoweritheta):

okay may be the assumption is that the events ar disjoint , then u can add them as normal arithmetic , and if their sum=1 they are complementary

OpenStudy (anonymous):

sorry I don't understand?

OpenStudy (anonymous):

oh never mind I got it. thanks for the help!